# JWST COMPASS: Insights into the Systematic Noise Properties of NIRSpec/G395H From a Uniform Reanalysis of Seven Transmission Spectra

### v1

@@ 2.2.2 White Light curve Fitting Procedure
For each target we fit the white lightcurves for each detector separately. We produce three fits for each target and detector using three different systematics models in order to investigate the impact of each model on the final white lightcurves. The first, which we refer to as the “0-vector” model, is a two-parameter model in which we fit for the out-of-transit flux and a linear slope in time. Fitting this slope has become standard for NIRSpec lightcurves in order to account for the pronounced downward trend in flux over time that is seen in NRS1, and the smaller slope seen in NRS2 (Espinoza2023). This 0-vector model provides a baseline against which we can compare the two more complex systematics models.
@@
The second model, which we refer to as the “shifts-only” model, is a four-parameter model defined by where and are the trace positions in the x and y-directions respectively recorded as offsets from the position of the integration-averaged trace position. Systematics models like this one, that account for variations in the telescope pointing over time, have been used in many recent JWST observations including NIRSpec/G395H observations (e.g. Rustamkulov2022; Teske2025; Alderson2023; Scarsdale2024).
@@
The third systematics model, which we refer to as the “6-vector” model, is an eight-parameter model consisting of the out-of-transit flux and linear slope from the 0-vector model plus coefficients for the first six principal components of the RPF timeseries described above. We choose to fit six components because we find that every observation has at most six vectors which display correlated noise. In the interest of carrying out a uniform analysis, we elect to use the same number of principal components for each observation, and using fewer than six would result in excluding principal components that show correlated noise for some observations, while using more than six would mean including components that display only white noise.
@@
We parameterize the transit in terms of the period, , planet/star radius ratio, , time of the transit relative to the start of the observation , semimajor axis in units of stellar radii , inclination , eccentricity and argument of periastron . Because the JWST observations do not meaningfully constrain , , , or , we place informative Gaussian priors on these parameters that are derived from the reported maximum likelihood values and uncertainties from the studies referenced in Table 3. We place uninformative uniform priors on (uniform between 0 and 1), (uniform between the beginning and end of the observation), and (uniform between where and are the mean and standard deviation of the reported value from the reference in 3) in order to allow the data to determine the posteriors for these parameters. For GJ 357 b and L 168-9 b the studies referenced in Table 3 assumed a fixed eccentricity of zero. For these cases we also set eccentricity to zero rather than fitting for its value. For the limb-darkening parameters we use informative priors computed with ExoTiC-LD (Grant2024) using the MPS-ATLAS-1 stellar models with the stellar parameters given in Table 5.
@@
| GJ 357 b | Oddo2023 | eccentricity set to 0 |
| TOI-836 b | Hawthorn2023 | |
| TOI-836 c | Hawthorn2023 | |
| TOI-776 b | Fridlund2024 | |
| TOI-776 c | Fridlund2024 | |
| L 98-59 c | Demangeon2021 | |
| L 168-9 b | Hobson2024 | eccentricity set to 0 |
@@
We use MCMC sampling to infer the parameters of the transit and systematics model. We run 60 chains for 100,000 samples each, discarding the first 5,000 samples of each chain as burn-in, for a total of 5,700,000 samples. We simultaneously model all available visits of each target, but we model the white light curves for the NRS1 and NRS2 detectors separately in order to account for differences in the properties of the detectors, and to assess the agreement in the transit parameters between them. In general, we find good agreement between the transit parameters for each detector with some observations showing statistically significant offsets in the transit time between NRS1 and NRS2 as has been seen in previous analysis of these targets (e.g. Wallack2024). Table 4 summarizes our posterior distributions for , , and . Because the , , , and are not well constrained by these observations, the posteriors are nearly identical to the priors and we elect not to report them in order to avoid confusion.
@@
In Figure 1 we plot 100 randomly drawn samples of the combined transit and systematics model from the posterior distribution for TOI-836 b, which we have chosen as a representative target. Corresponding plots for the full set of targets are available as a figure set in the HTML version of this article. Note that, while the systematics model primarily fits short timescale variability, and does not capture longer timescale variations in flux such as the pre-transit bump in NRS1. In Section 3 we discuss in detail the variability timescales present in the principle component vectors and explore the origins of this variability. Figure 2 shows the posterior distribution for the transit parameters for the white light curve of TOI-836 b. Corner plots for the full set of targets are also available as a figure set in the HTML version of the article. Figure 3 shows the improvement in the degree of red noise in the residuals when using the 6-vector systematics model (colored lines) in comparison to a 0-vector model (gray lines, left side) and the shifts-only model (gray lines, right side). From these plots we can see that use of the 6-vector model results in less red noise than either of the alternative models in the majority of cases. The difference is particularly notable for the NRS1 detector and for the three-group observations (GJ 357 b, TOI-836 b, and TOI-836 c). The improvement is less significant for the NRS2 detector and for the four-group and seven-group observations (L 98-59 c, L 168-9 b, TOI-776 b and TOI-776 c), though, for unknown reasons, the second visit of TOI-776 b does show significant improvement in the red noise for the NRS1 white lightcurve.
@@
In Figures 4 and 5 we plot the difference in the Bayesian Information Criterion () and the Akaike Information Criterion () between the 6-vector model and the 0-vector model, and between the 6-vector model and the shifts-only model. For each comparison a higher value of or indicates a preference for the 6-vector model. Following Raftery1995, we consider a or to indicate strong evidence for the 6-vector model and or to indicate strong evidence against the 6-vector model, as indicated by the shaded region in the figures. This corresponds roughly to an odds ratio of 150:1 in favor of the selected hypothesis. The difference between the Akaike and Bayesian Information Criterion is in how they penalize model complexity: the Bayesian Information Criterion is smaller by for each additional model parameter, where is the number of samples in the dataset (in this case the number of integrations in the timeseries). In contrast, the Akaike Information Criterion is reduced only by for each additional model parameter. The result is that the BIC appears less favorable for the more complex 6-vector model compared to the 0-vector model than does the AIC.
2.3 Spectral Light curve Analysis
We independently fit the light curves for each spectral bandpass, holding all transit parameters constant except for the planetary radius and the two quadratic limb-darkening parameters. The planetary radius, out-of-transit flux level, slope of the linear trend, and PCA vector coefficients are unconstrained. We fit the spectral light curves in 0.02 micron wide bins, corresponding to approximately 30 pixels-per-bin or a spectral resolution of R=200 at 4 m, by running 10,000 steps of MCMC, discarding the first 5,000 as burn-in for each of 36 chains to produce posterior distributions for all parameters.
| planet name | mass (M⊕) | Teq (K) | radius (R⊕) | Teff | R () | reference | ||
| L 168-9 b | 4.07 | 998 | 1.58 | 3842 | 0.60 | 4.84 | -0.03 | Hobson2024 |
| L 98-59 c | 2.22 | 553 | 1.39 | 3415 | 0.30 | 4.45 | -0.46 | Demangeon2021 |
| GJ 357 b | 1.84a | 525 | 1.33 | 3505 | 0.34 | 4.94 | -0.12 | Oddo2023 |
| TOI-776 c | 6.90 | 420 | 2.02 | 3725 | 0.55 | 4.80 | -0.21 | Fridlund2024 |
| TOI-776 b | 5.00 | 520 | 1.82 | 3725 | 0.55 | 4.80 | -0.21 | Fridlund2024 |
| TOI-836 b | 4.53 | 871 | 1.79 | 4552 | 0.67 | 4.74 | -0.28 | Hawthorn2023 |
| TOI-836 c | 9.60 | 665 | 2.50 | 4552 | 0.67 | 4.74 | -0.28 | Hawthorn2023 |
(a) Luque2019
We produce two transmission spectra for each planet: one with the same 6-vector systematics model used for the white light curves and one without the basis vectors that includes only the out-of-transit flux and the slope of the linear trend (the 0-vector model). Figure 7 shows both versions of the transmission spectra along with the residuals between the two, and figure 8 overplots the precision for both versions of the transmission spectra along with the pandexo predicted precision. In Table 6 we give approximate inflation factors to recover the observed precision given the pandexo prediction for each version of the spectra. We also experimented with optimizing the number of basis vectors in the systematics model for each spectral bandpass by testing every possible subset of the six basis vectors using a generalized least-squares fit and finding the subset that minimized the reduced . We found that there was no subset of the six basis vectors for which the fitted model had a substantially improved over the full 6-vector model, and we further found that the optimized systematics models resulted in spectra that were nearly identical to the 6-vector systematics model for every planet.
@@ 2.3 Spectral Light curve Analysis
We found that the difference between the spectra obtained with each systematics model was only significant for three planets: GJ 357 b, TOI-836 b, and TOI-836 c. These targets are notable in that they were observed with three groups, while the remaining four targets were all observed with four or more groups. In Section 3 we investigate the appearance of the systematics as a function of group number and demonstrate that the systematics are stronger for lower group numbers. For the other four planets in the sample, we determined that the use of the more complex 6-vector model for the spectral light curve fitting was not justified since it did not alter the final spectrum. We therefore elected to use the transmission spectra for the 6-vector systematics model for these three planets and the 0-vector systematics model for the remaining planets. The final versions of all spectra are shown in Figure 6.
3 NIRSpec Detector Systematics
@@
Following this same logic, the second principal component must be associated with the rotation of the trace. To the left of the pixel column this component is tracking an upward motion in the trace, and to the right of the pixel column this component tracks a downward shift. The third and fourth PCA vector both appear to originate from a periodic de-focusing of the trace, as evidenced by the switch in the sign of the correlation as we move from the edge of the trace to its center. The correlation map for the fourth PCA vector shows a similar structure, except that the sign of the correlation changes twice as we move from the edge of the trace to the center between the 500th and 1000th pixel columns, suggesting that this is related to a higher-order change in the shape of the trace’s cross section. The map for the fifth component shows faint horizontal striping, pointing towards residual 1/f noise (this striping appears in the fourth and 6th components as well, indicating that the correspondence between the source of some variability and the principal components is not one-to-one). Both the fifth and sixth components show pattern of alternating positively and negatively correlated pixel rows, suggesting that this signal is related to alternating column noise (ACN) (Rauscher2015). Note that the engineering and science reference frames are rotated ninety degrees relative to each other, so that the “columns” referenced in “alternating column noise” are actually seen as rows in the science data.
@@
The full set of components identified above does not appear in every single visit and detector. The most common components are the shifting component, which appears universally, and the de-focusing, rotation, and ACN components, which appear in the majority of observations. In addition to these common components, some visits and detectors display additional variability that typically arises from a small subset of pixels that do not show oscillatory behavior. Examples of these components are the third and sixth components from the NRS2 detector in Figure 9. The origin of these systematics is highly uncertain, but the discontinuities seen in the sixth component of NRS2 may be related to the “popcorn mesa” noise studied by Rauscher2004, in which individual pixels appeared to switch between high and low voltage states during testing of a prototype of the NIRSpec detectors. In Figure 9 the white squares mark the locations of pixels highly correlated with the corresponding principal component. In most cases these pixels do not coincide with the trace and therefore are likely to have a negligible effect on the extracted light curves. It may be possible to identify and remove these pixels prior to spectral extraction, however this is a topic that we leave to future work.
@@
A strong candidate for the oscillations on 3.4 and 3.63-minute timescales is thermal cycling of the heaters in the ISIM Electronics Compartment, which according to Rigby2024 results in “a semi-periodic variation primarily in astigmatism” on timescales of 2-4 minutes. This corresponds well to the observed timescales of the oscillations and explains their prominence in the de-focusing component, which would correspond to the astigmatism discussed in Rigby2024. Why these oscillation timescales also appear in the rotation component is unclear, as is the origin of the rotation itself.
@@
3.1 Dependence on Group Number
@@
For NRS1, both the average amplitude of each component and the scatter in the amplitudes between targets and visits decreases as we move from three groups to four groups. The amplitude of the shifting and de-focusing components decreases further as we move to seven groups, while it appears unchanged for the ACN and rotating components between four and seven groups. These results confirm that low group number observations should be treated with caution, as has been noted in previous studies of datasets with small group numbers (Alderson2024; Sikora2025). Given that the de-focusing component reflects systematics on 2-4 minute timescales, this may be especially important for analyses that depend on the shape of the transit ingress and egress, such as transit timing/duration variations, limb-darkening measurements, investigations into planetary and/or stellar oblateness.
Recently, Luque2024 applied a similar ICA-based systematics model to phase curve observations of TOI-1685 b taken with 16 groups and found that the model was unsuccessful at removing correlated noise. Their Figures 7 and 8 show their ICA basis vectors and associated Lomb-Scargle periodograms, which do not exhibit the same periodic behavior we observe in our PCA basis vectors. This additional datapoint supports the notion that systematics related to time-variable trace morphologies are less of any issue for observations using high numbers of groups per integration.
3.2 Wavelength Dependence

### v2

@@ II.2.2 White Light curve Fitting Procedure
For each target we fit the white lightcurves for each detector separately. We produce three fits for each target and detector using three different systematics models in order to investigate the impact of each model on the final white lightcurves. The first, which we refer to as the “0-vector” model, is a two-parameter model in which we fit for the out-of-transit flux and a linear slope in time. Fitting this slope has become standard for NIRSpec lightcurves in order to account for the pronounced downward trend in flux over time that is seen in NRS1, and the smaller slope seen in NRS2 (Espinoza et al., 2023). This 0-vector model provides a baseline against which we can compare the two more complex systematics models.
@@
The second model, which we refer to as the “shifts-only” model, is a four-parameter model defined by where and are the trace positions in the x and y-directions respectively recorded as offsets from the position of the integration-averaged trace position. Systematics models like this one, that account for variations in the telescope pointing over time, have been used in many recent JWST observations including NIRSpec/G395H observations (e.g. Rustamkulov et al., 2022; Teske et al., 2025; Alderson et al., 2023; Scarsdale et al., 2024).
@@
The third systematics model, which we refer to as the “6-vector” model, is an eight-parameter model consisting of the out-of-transit flux and linear slope from the 0-vector model plus coefficients for the first six principal components of the RPF timeseries described above. We choose to fit six components because we find that every observation has at most six components which display correlated noise, with higher order components appearing consistent with white noise when inspected by eye. In the interest of carrying out a uniform analysis, we elect to use the same number of principal components for each observation, and using fewer than six would result in excluding principal components that show correlated noise for some observations, while using more than six would mean including components that display only white noise.
@@
We parameterize the transit in terms of the period, , planet/star radius ratio, , time of the transit relative to the start of the observation , semimajor axis in units of stellar radii , inclination , eccentricity and argument of periastron . Because the JWST observations do not meaningfully constrain , , , or , we place informative Gaussian priors on these parameters that are derived from the reported maximum likelihood values and uncertainties from the studies referenced in Table 3. We place uninformative uniform priors on (uniform between 0 and 1), (uniform between the beginning and end of the observation), and (uniform between where and are the mean and standard deviation of the reported value from the reference in 3) in order to allow the data to determine the posteriors for these parameters. For GJ 357 b and L 168-9 b the studies referenced in Table 3 assumed a fixed eccentricity of zero. For these cases we also set eccentricity to zero rather than fitting for its value. For the limb-darkening parameters we use informative priors computed with ExoTiC-LD (Grant and Wakeford, 2024) using the MPS-ATLAS-1 stellar models with the stellar parameters given in Table 4.
@@
| GJ 357 b | Oddo et al. (2023) | eccentricity set to 0 |
| TOI-836 b | Hawthorn et al. (2023) | |
| TOI-836 c | Hawthorn et al. (2023) | |
| TOI-776 b | Fridlund et al. (2024) | |
| TOI-776 c | Fridlund et al. (2024) | |
| L 98-59 c | Demangeon et al. (2021) | |
| L 168-9 b | Hobson et al. (2024) | eccentricity set to 0 |
@@
We use MCMC sampling to infer the parameters of the transit and systematics model. We run 60 chains for 100,000 samples each, discarding the first 5,000 samples of each chain as burn-in, for a total of 5,700,000 samples. We simultaneously model all available visits of each target, but we model the white light curves for the NRS1 and NRS2 detectors separately in order to account for differences in the properties of the detectors, and to assess the agreement in the transit parameters between them. We obtain effective samples for all parameters and all targets. In general, we find good agreement between the transit parameters for each detector with some observations showing statistically significant offsets in the transit time between NRS1 and NRS2 as has been seen in previous analysis of these targets (e.g. Wallack et al., 2024). Table 5 summarizes our posterior distributions for , , and . Because the , , , and are not well constrained by these observations, the posteriors are nearly identical to the priors and we elect not to report them in order to avoid confusion.
@@
In Figure 1 we plot 100 randomly drawn samples of the combined transit and systematics model from the posterior distribution for TOI-836 b, which we have chosen as a representative target. Corresponding plots for the full set of targets are available as a figure set in the HTML version of this article. Note that, while the systematics model primarily fits short timescale variability, and does not capture longer timescale variations in flux such as the pre-transit bump in NRS1. In Section III we discuss in detail the variability timescales present in the principle component vectors and explore the origins of this variability. Figure 2 shows the posterior distribution for the transit parameters for the white light curve of TOI-836 b. Corner plots for the full set of targets are also available as a figure set in the HTML version of the article. Figure 3 shows the improvement in the degree of red noise in the residuals when using the 6-vector systematics model (colored lines) in comparison to a 0-vector model (gray lines, left side) and the shifts-only model (gray lines, right side). From these plots we can see that use of the 6-vector model results in less red noise than either of the alternative models in the majority of cases. The difference is particularly notable for the NRS1 detector and for the three-group observations (GJ 357 b, TOI-836 b, and TOI-836 c). The improvement is less significant for the NRS2 detector and for the four-group and seven-group observations (L 98-59 c, L 168-9 b, TOI-776 b and TOI-776 c), though, for unknown reasons, the second visit of TOI-776 b does show significant improvement in the red noise for the NRS1 white lightcurve.
| planet name | mass (M⊕) | Teq (K) | radius (R⊕) | Teff | R () | reference | ||
| L 168-9 b | 4.07 | 998 | 1.58 | 3842 | 0.60 | 4.84 | -0.03 | Hobson et al. (2024) |
| L 98-59 c | 2.22 | 553 | 1.39 | 3415 | 0.30 | 4.45 | -0.46 | Demangeon et al. (2021) |
| GJ 357 b | 1.84a | 525 | 1.33 | 3505 | 0.34 | 4.94 | -0.12 | Oddo et al. (2023) |
| TOI-776 c | 6.90 | 420 | 2.02 | 3725 | 0.55 | 4.80 | -0.21 | Fridlund et al. (2024) |
| TOI-776 b | 5.00 | 520 | 1.82 | 3725 | 0.55 | 4.80 | -0.21 | Fridlund et al. (2024) |
| TOI-836 b | 4.53 | 871 | 1.79 | 4552 | 0.67 | 4.74 | -0.28 | Hawthorn et al. (2023) |
| TOI-836 c | 9.60 | 665 | 2.50 | 4552 | 0.67 | 4.74 | -0.28 | Hawthorn et al. (2023) |
(a) Luque et al. (2019)
In Figures 4 and 5 we plot the difference in the Bayesian Information Criterion () and the Akaike Information Criterion () between the 6-vector model and the 0-vector model, and between the 6-vector model and the shifts-only model. For each comparison a higher value of or indicates a preference for the 6-vector model. Following Raftery (1995), we consider a or to indicate strong evidence for the 6-vector model and or to indicate strong evidence against the 6-vector model, as indicated by the shaded region in the figures. This corresponds roughly to an odds ratio of 150:1 in favor of the selected hypothesis. The difference between the Akaike and Bayesian Information Criterion is in how they penalize model complexity: the Bayesian Information Criterion is smaller by for each additional model parameter, where is the number of samples in the dataset (in this case the number of integrations in the timeseries). In contrast, the Akaike Information Criterion is reduced only by for each additional model parameter. The result is that the BIC appears less favorable for the more complex 6-vector model compared to the 0-vector model than does the AIC.
@@
II.3 Spectral Light curve Analysis
We independently fit the light curves for each spectral bandpass, holding all transit parameters constant at the maximum-likelihood values from the white lightcurve posteriors except for the planetary radius and the two quadratic limb-darkening parameters. The planetary radius, out-of-transit flux level, slope of the linear trend, and PCA vector coefficients are unconstrained. We fit the spectral light curves in 0.02 micron wide bins, corresponding to approximately 30 pixels-per-bin or a spectral resolution of R=200 at 4 m, by running 10,000 steps of MCMC, discarding the first 5,000 as burn-in for each of 36 chains to produce posterior distributions for all parameters.
We produce two transmission spectra for each planet: one with the same 6-vector systematics model used for the white light curves and one without the basis vectors that includes only the out-of-transit flux and the slope of the linear trend (the 0-vector model). Figure 7 shows both versions of the transmission spectra along with the residuals between the two, and Figure 8 overplots the precision for both versions of the transmission spectra along with the pandexo predicted precision. In Table 6 we give approximate inflation factors to recover the observed precision for the transit depth given the pandexo prediction for the transit depth uncertainty for each version of the spectra.
We also experimented with optimizing the number of basis vectors in the systematics model for each spectral bandpass by testing every possible subset of the six basis vectors using a generalized least-squares fit and finding the subset that minimized the reduced . We found that there was no subset of the six basis vectors for which the fitted model had a substantially improved reduced over the full 6-vector model, and we further found that the optimized systematics models resulted in spectra that were nearly identical to the 6-vector systematics model for every planet.
@@ II.3 Spectral Light curve Analysis
We found that the difference between the spectra obtained with each systematics model was only significant for three planets: GJ 357 b, TOI-836 b, and TOI-836 c. These targets are notable in that they were observed with three groups, while the remaining four targets were all observed with four or more groups. In Section III we investigate the appearance of the systematics as a function of group number and demonstrate that the systematics are stronger for lower group numbers. For the other four planets in the sample, we determined that the use of the more complex 6-vector model for the spectral light curve fitting was not justified since it did not alter the final spectrum. We therefore elected to use the transmission spectra for the 6-vector systematics model for these three planets and the 0-vector systematics model for the remaining planets. The final versions of all spectra are shown in Figure 6.
III NIRSpec Detector Systematics
@@
Following this same logic, the second principal component must be associated with the rotation of the trace. To the left of the pixel column this component is tracking an upward motion in the trace, and to the right of the pixel column this component tracks a downward shift. The third and fourth PCA vector both appear to originate from a periodic de-focusing of the trace, as evidenced by the switch in the sign of the correlation as we move from the edge of the trace to its center. The correlation map for the fourth PCA vector shows a similar structure, except that the sign of the correlation changes twice as we move from the edge of the trace to the center between the 500th and 1000th pixel columns, suggesting that this is related to a higher-order change in the shape of the trace’s cross section. The map for the fifth component shows faint vertical striping, pointing towards residual 1/f noise (this striping appears in the fourth and 6th components as well, indicating that the correspondence between the source of some variability and the principal components is not one-to-one). Both the fifth and sixth components show pattern of alternating positively and negatively correlated pixel rows, suggesting that this signal is related to alternating column noise (ACN) (Rauscher, 2015). Note that the engineering and science reference frames are rotated ninety degrees relative to each other, so that the “columns” referenced in “alternating column noise” are actually seen as rows in the science data.
@@
The full set of components identified above does not appear in every single visit and detector. The most common components are the shifting component, which appears universally, and the de-focusing, rotation, and ACN components, which appear in the majority of observations. In addition to these common components, some visits and detectors display additional variability that typically arises from a small subset of pixels that do not show oscillatory behavior. Examples of these components are the third and sixth components from the NRS2 detector in Figure 9. The origin of these systematics is highly uncertain, but the discontinuities seen in the sixth component of NRS2 may be related to the “popcorn mesa” noise studied by Rauscher et al. (2004), in which individual pixels appeared to switch between high and low voltage states during testing of a prototype of the NIRSpec detectors. In Figure 9 the white squares mark the locations of pixels highly correlated with the corresponding principal component. In most cases these pixels do not coincide with the trace and therefore are likely to have a negligible effect on the extracted light curves. It may be possible to identify and remove these pixels prior to spectral extraction, however this is a topic that we leave to future work.
@@
A strong candidate for the oscillations on 3.4 and 3.63-minute timescales is thermal cycling of the heaters in the ISIM Electronics Compartment, which according to Rigby et al. (2024) results in “a semi-periodic variation primarily in astigmatism” on timescales of 2-4 minutes. This corresponds well to the observed timescales of the oscillations and explains their prominence in the de-focusing component, which would correspond to the astigmatism discussed in Rigby et al. (2024). Why these oscillation timescales also appear in the rotation component is unclear, as is the origin of the rotation itself.
@@
III.1 Dependence on Group Number
@@
For NRS1, both the average amplitude of each component and the scatter in the amplitudes between targets and visits decreases as we move from three groups to four groups. The amplitude of the shifting and de-focusing components decreases further as we move to seven groups, while it appears unchanged for the ACN and rotating components between four and seven groups. These results confirm that low group number observations should be treated with caution, as has been noted in previous studies of datasets with small group numbers (Alderson et al., 2024; Sikora et al., 2025). Given that the de-focusing component reflects systematics on 2-4 minute timescales, this may be especially important for analyses that depend on the shape of the transit ingress and egress, such as transit timing/duration variations, limb-darkening measurements, investigations into planetary and/or stellar oblateness.
Recently, Luque et al. (2024) applied a similar ICA-based systematics model to phase curve observations of TOI-1685 b taken with 16 groups and found that the model was unsuccessful at removing correlated noise. Their Figures 7 and 8 show their ICA basis vectors and associated Lomb-Scargle periodograms, which do not exhibit the same periodic behavior we observe in our PCA basis vectors. This additional datapoint supports the notion that systematics related to time-variable trace morphologies are less of any issue for observations using high numbers of groups per integration.
III.2 Wavelength Dependence